OpenStudy (anonymous):

Simplify x 11th power ------------ a -2 power

OpenStudy (lgbasallote):

this question is a little vague for me... is it \[\huge \frac{x^{11}}{a^{-2}}\]

OpenStudy (anonymous):

@lgbasallote Yes

OpenStudy (anonymous):

@lgbasallote oops! Its "a" not "x" for the top one

OpenStudy (lgbasallote):

ahh so \[\huge \frac{a^{11}}{a^{-2}}\]

OpenStudy (anonymous):

wait no! the bottom exponent is 7. That's the correct problem

OpenStudy (lgbasallote):

...so \[\huge \frac{a^{11}}{a^{-7}}\]??

OpenStudy (anonymous):

no negative sign

OpenStudy (lgbasallote):

so \[\huge \frac{a^{11}}{a^7}\]

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

are you familiar with the algebraic rule \[\huge \frac{x^m}{x^n} \implies x^{m-n}\]

OpenStudy (anonymous):

Yes but for some reason i thought that the problem was division, not a fraction. I'm not really sure which one it is

OpenStudy (lgbasallote):

fraction is division

OpenStudy (anonymous):

Oh.

OpenStudy (lgbasallote):

so do you know how to solve \(\large \frac{a^{11}}{a^7}\) now?

OpenStudy (anonymous):

is it 2a to the 4th power? ---> 2a4

OpenStudy (lgbasallote):

where did 2 come from?

OpenStudy (anonymous):

because there's 2 "a's"

OpenStudy (lgbasallote):

not exactly... like i said before \[\Large \frac{x^m}{x^n} \implies x^{m-n}\] so \[\Large \frac{a^{11}}{a^7} \implies a^{11 - 7} \implies a^4\] no 2

OpenStudy (anonymous):

Oh! ok! Thank you

OpenStudy (lgbasallote):

welcome